When 1.0 m3 of an ideal gas expands to 2.5 m3 at a constant pressure of 1000 N/m2, how much of work is done by the gas in J?

1
Expert's answer
2021-11-24T12:01:25-0500

"W=p(V_2-V_1)=1000\cdot(2.5-1)=1500\ (J)" . Answer